The following data were collected on the height (inches) and weight (pounds) of women swimmers where height is the independent variable and weight is the dependent variable.

Height (x) = 68, 64, 62, 65, 66
Weight (y) = 132, 108, 102, 115, 128

a)Compute SSE, SST, and SSR

b)Compute the coefficient of determination r2. Comment on the goodness of fit.

c)Compute the sample correlation coefficient.
Answer #1

a)

SSE =Syy-(Sxy)2/Sxx= 51.000
SST=Syy= 656.0000
SSR =(Sxy)2/Sxx = 605.0000

b)

Coeffficient of determination R^2 =SSR/SST= 0.9223

since R2 is very high, this tells us that values fit in the model very well

c)

correlation coefficient r= Sxy/(√Sxx*Syy) = 0.9603